Reports 1-1 of 1 Clear search Modify search
VIS (General)
satoru.ikeda - 12:04 Friday 06 August 2021 (17831) Print this report
Routine confirmation of DGS maintenance day

Routine confirmation for DGS maintenance day.

Stepper motor
 [Stepper motor]
  Checked if the power supply can be controlled by BIO.
  Checked if the script can be started.
  Confirmed that the limit switch installed is green.
 [Check result]
 ETMY:
  Power control: No problem
  Limit SW: currently all NG, but will be adjusted

 ITMY:
  Power control: No problem
  Limit SW: F0YAW is NG (known)

 Other than above: No problem

Release Sensor
 [Confirmation]
 Confirmed that Release Sensor is 0. 1: Seated
 [Check result]
 ETMY, ITMX, ITMY : No problem
 ETMX_F1_1 : Abnormal, others are OK
 

Images attached to this report
Non-image files attached to this report
Comments to this report:
satoru.ikeda - 13:19 Friday 20 August 2021 (17968) Print this report

Routine confirmation of DGS maintenance day(Aug.20)

Performed routine check of DGS maintenance day.
No problems other than known problems.

Stepper motor
 [Checked]
  Checked if the power supply can be controlled by BIO.
  Checked if Script can be started.
  Check that the limit switch installed is green.
 [Check result]
 ETMY:
  Power control: No problem
  Limit SW: currently all NG, but will be adjusted in the future (known)

 ITMY:
  Power control: No problem
  Limit SW: F0YAW is NG (known)

 Other than above: No problem

Release Sensor
 [Confirmation]
 Confirmed that Release Sensor is 0. 1: Seated
 [Check result]
 ETMY, ITMX, ITMY : No problem
 ETMX_F1_1 : Abnormal, no other problems (known)
 

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 14:09 Friday 27 August 2021 (18040) Print this report

Routine confirmation of DGS maintenance day(Aug.27)

Performed routine check of DGS maintenance day.
No problems other than known problems.

Stepper motor
 [Checked]
  Checked if the power supply can be controlled by BIO.
  Checked if Script can be started.
  Check that the limit switch installed is green.
 [Check result]
 ETMY:
  Power control: No problem
  Limit SW: currently all NG, but will be adjusted in the future (known)

 ITMY:
  Power control: No problem
  Limit SW: F0YAW is NG (known)

 Other than above: No problem

Release Sensor
 [Confirmation]
 Confirmed that Release Sensor is 0. 1: Seated
 [Check result]
 ETMY, ITMX, ITMY : No problem
 ETMX_F1_1 : Abnormal, no other problems (known)
 

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 17:23 Friday 03 September 2021 (18144) Print this report

Routine confirmation of DGS maintenance day(Sep.3)

Performed routine check of DGS maintenance day.
No problems other than known problems.

Stepper motor
 [Checked]
  Checked if the power supply can be controlled by BIO.
  Checked if Script can be started.
  Check that the limit switch installed is green.
 [Check result]
 ETMY:
  Power control: No problem
  Limit SW: currently all NG, but will be adjusted in the future (known)

 ITMY:
  Power control: No problem
  Limit SW: F0YAW is NG (known)

 Other than above: No problem

Release Sensor
 [Confirmation]
 Confirmed that Release Sensor is 0. 1: Seated
 [Check result]
 ETMY, ITMX, ITMY : No problem
 ETMX_F1_1 : Abnormal, no other problems (known)
 

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 14:10 Friday 10 September 2021 (18227) Print this report

Routine confirmation of DGS maintenance day (Sep.10)

Routine checks for DGS maintenance day (Sep.10) were performed.
There are no problems except for the known ones.

Stepper motor
 [Checked that the power supply can be controlled by BIO.]
  Checked if power can be controlled by BIO.
  Check if Script can be started.
  Check that the limit switch installed is green.
 [Check result]
  No problem

Release Sensor
 [Checked]
 Confirmed that the Release Sensor is 0. 1: Seated
 [Check result]
 ETMY, ITMX, ITMY : No problem
 ETMX_F1_1 : abnormal, no other problems (known)

Picomotor
 [Checked]
  PingTest is performed.
 [Result]

 NG:
  ETMX: NG
  SR2(IM): NG
   => Rebooted and recovered.
  MCI,STM1,IMMT1,IMMT2:NG
  AF_WFS,POP_POM,REFL_WFS,POP1,STM2: NG
   => Confirmation only, no recovery work done.
 OK:
  ITMX,ITMY,ETMY
  BS(IM,BF),SRM(IM,BF),SR2(BF),SR3(IM,BF)
  PRM(IM,BF),PR2(IM,BF),PR3(IM,BF)
  MCF(MCO,MCI),MCE
  OMMT1,OMMT2,OSTM
  PCAL(EX1,EX2,EY1,EY2):Not confirmed because it is ON only when used.

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 15:57 Friday 17 September 2021 (18295) Print this report

Routine confirmation of DGS maintenance day (9/17)

Routine checks for DGS maintenance day (9/17) were performed.
There are no problems except for the known ones.

Stepper motor
 [Checked that the power supply can be controlled by BIO.]
  Checked if power can be controlled by BIO.
  Check if Script can be started.
  Check that the limit switch installed is green.
 [Check result]
  No problem

Release Sensor
 [Checked]
 Confirmed that the Release Sensor is 0. 1: Seated
 [Check result]
 ETMY, ITMX, ITMY : No problem
 ETMX_F1_1 : abnormal, no other problems (known)

Picomotor
 [Checked]
  PingTest is performed.
 [Result]

 NG:
  BS_BF:NG
   => Rebooted and recovered.
  SR2_STM:NG
   => The power supply was almost disconnected, so I asked Hirata-san to plug it back in.
  MCI,STM1,IMMT1,IMMT2:NG
  AS_WFS,POP_POM,POS_POM,REFL_WFS,POM1: NG
   => Confirmation only, no recovery work done.

 OK:
  ITMX,ITMY,ETMX,ETMY
  BS(IM,BF),SRM(IM,BF),SR2(IM,BF,STM),SR3(IM,BF)
  PRM(IM,BF),PR2(IM,BF),PR3(IM,BF)
  MCF(MCO,MCI),MCE
  OMMT1,OMMT2,OSTM

  PCAL(EX1,EX2,EY1,EY2):Not confirmed because it is ON only when used.
  

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 17:13 Friday 24 September 2021 (18347) Print this report

Routine confirmation of DGS maintenance day (9/24)

Routine checks for DGS maintenance day (9/24) were performed.
There are no problems except for the known ones.

Stepper motor
 [Checked that the power supply can be controlled by BIO.]
  Checked if power can be controlled by BIO.
  Check if Script can be started.
  Check that the limit switch installed is green.
 [Check result]
  No problem

Release Sensor
 [Checked]
 Confirmed that the Release Sensor is 0. 1: Seated
 [Check result]
 ETMY, ITMX, ITMY : No problem
 ETMX_F1_1 : abnormal, no other problems (known)

Picomotor
 [Checked]
  PingTest is performed.
 [Result]

 NG:
  SR3_IM: NG
  ETMX: NG
   ETMY: NG => Both rebooted and came back.
  ETMY: NG
   => Can't connect to the rebooter either. Power disconnected? Unknown cause. Will check on Monday.
  SR2_STM: NG
   => The rebooter is not installed yet, so I will install it next week and get it back.

  MCI,STM1,IMMT1,IMMT2:NG
  AS_WFS,POP_POM,POS_POM,REFL_WFS,POM1: NG
   => Confirmation only, no recovery work done.

 OK:
  ITMX,ITMY
  BS(IM,BF),SRM(IM,BF),SR2(IM,BF),SR3(BF)
  PRM(IM,BF),PR2(IM,BF),PR3(IM,BF)
  MCF(MCO,MCI),MCE
  OMMT1,OMMT2,OSTM

  PCAL(EX1,EX2,EY1,EY2):Not confirmed because it is ON only when used.
 

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 16:13 Monday 27 September 2021 (18356) Print this report

[Ikeda, Yasui-san, Nakagaki-san]

Two issues found during routine maintenance have been addressed.

ETMY: NG -> OK
 => Can't connect to the rebooter either. Was the power supply disconnected? Cause unknown. Will check on Monday. 
 
 [Investigation result]
  It was found that Picomotor and Picomotor's rebooter did not respond to ping during the regular maintenance on 9/27.
  In the subsequent investigation, it was found that the temperature control device did not respond to pings as well.
  When we checked the log of the temperature control device, we found that the connection was lost around 11:23 on September 20.
  When I checked the log of the temperature control device, I found that the connection had been lost around 11:23 a.m. on 9/20. The backbone LAN cable was plugged into the power supply port of the network SW (S2112636:GS108TV3) placed in the VIS mini rack.
   This LED was turned off and it was found to be unconnected.
 This network SW had an AC adapter and a LAN connected to the PoE power supply port.
   The network was restored by removing the AC adapter or connecting the AC adapter to a port other than the power supply port.
 [Details]
  The AC adapter was removed because it was found that PoE power supply was available.
   Confirmed that it is connected to the network.

SR2_STM: NG -> OK
 => The rebooter is not installed yet, so I will install it next week to get it back.
 => Installed a switch (S2113327) for power control in Picomotor of SR2_STM.

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 14:18 Friday 01 October 2021 (18418) Print this report

Routine confirmation of DGS maintenance day (10/01)

Routine check for DGS maintenance day (10/01) was performed.
Two problems were found.
The limit switch of BS_F0YAW was in contact with both L and R.

The LAN of ETMY Pico and the temperature control device could not be connected.
 => I changed the power supply of the network SW to PoE on 9/27, but it recurred after that.

Stepper motor
 [Checked]
  Checked if the power can be controlled by BIO.
  Checked if Script can be started.
  Checked that the limit switch installed is green.
 [Check result]
  The limit switch of BS_F0YAW was abnormal because both L and R were in contact.
  => Confirmation required.
  No other problems.

Release Sensor
 [Confirmation]
 Confirmed that Release Sensor is 0. 1: Seated
 [Check result]
 ETMY, ITMX, ITMY : No problem
 ETMX_F1_1 : abnormal, no other problems (known)

Picomotor
 [Checked]
  PingTest is performed.
 [Result]
 NG:
  ETMY
   => Unable to access from LAN after 9/27 11:38.
    Will check on Monday.
  SR2_IM,SR2_STM
   => Both have been rebooted and recovered.

  MCI,STM1,IMMT1,IMMT2:NG
  AS_WFS,POP_POM,POS_POM,REFL_WFS,POM1: NG
   => Confirmation only, no recovery work done.

 OK:
  ETMX,ITMX,ITMY
  BS(IM,BF),SRM(IM,BF),SR2(BF),SR3(IM,BF)
  PRM(IM,BF),PR2(IM,BF),PR3(IM,BF)
  MCF(MCO,MCI),MCE
  OMMT1,OMMT2,OSTM

  PCAL(EX1,EX2,EY1,EY2):Not confirmed because it is ON only when used.

Temperature Control
 [Checked]
 Check if the temperature acquisition is working properly.

 [Check result]
 ETMY:NG
   => Unable to access from LAN after 9/27 11:38.
    Will check on Monday.
  Other than the above, everything was OK.
 

Images attached to this comment
Non-image files attached to this comment
naoatsu.hirata - 17:18 Tuesday 05 October 2021 (18459) Print this report

>Stepper motor

>[Check result]
> The limit switch of BS_F0YAW was abnormal because both L and R were in contact.
> => Confirmation required.

I connected the in-air cable, and confirmed limit switches work well.

Images attached to this comment
satoru.ikeda - 16:27 Wednesday 06 October 2021 (18479) Print this report

[Ikeda, Oshino-san].

About the network SW installed in EYV's VIS mini rack being turned off.
I asked Oshino-san to turn off and on the PoE power supply port (Port 17) to which the network SW of the VIS mini rack is connected.
(This was done at around 10:53 on 2021-09-27)
We decided to wait and see if the system works stably after the reboot.
Currently, as of 16:22, it is normal

Non-image files attached to this comment
satoru.ikeda - 16:28 Friday 08 October 2021 (18514) Print this report

Routine confirmation of DGS maintenance day (Oct.8)

Stepper motor
 [Contents of check]
  Checked if the power supply can be controlled by BIO.
  Check if Script can be started.
  Checked that the limit switch installed is in Green.
 [Check result]
  No problem

Release Sensor
 [Checked]
 Confirmed that the Release Sensor is 0. 1: Seated
 [Check result]
 ETMY, ITMX, ITMY : No problem
 ETMX_F1_1 : abnormal, no other problems (known)

Picomotor
 [Checked]
  PingTest is performed.
 [Result]
 NG:
  No problem

  MCI,STM1,IMMT1,IMMT2: NG
  AS_WFS,POP_POM,POS_POM,REFL_WFS,POM1: NG
   => Confirmation only, no recovery work done.

 OK:
  ETMX,ETMY,ITMX,ITMY
  BS(IM,BF),SRM(IM,BF),SR2(IM,BF,STM),SR3(IM,BF)
  PRM(IM,BF),PR2(IM,BF),PR3(IM,BF)
  MCF(MCO,MCI),MCE
  OMMT1,OMMT2,OSTM

  PCAL(EX1,EX2,EY1,EY2):Not confirmed because it is ON only when used.

Temperature regulator
 [Checked]
 Check if the temperature acquisition is working properly.

 [Check result]
  No problem
 

Images attached to this comment
Non-image files attached to this comment
satoru.ikeda - 16:51 Friday 15 October 2021 (18596) Print this report

Routine check on DGS maintenance day (Oct. 15)

Stepper motor
 [Checked]
  Check if the power supply can be controlled by BIO.
  Check if the script can be started.
  Confirmed that the limit switch installed is in Green.
 [Check result]
  No problem

Release Sensor
 [Checked]
 Confirmed that the Release Sensor is 0. 1: Seated
 [Check result]
 ETMY, ITMX, ITMY : No problem
 ETMX_F1_1 : abnormal, no other problems (known)

Picomotor
 [Checked]
  PingTest is performed.
 [Result]
 NG:
  No problem

  MCI,STM1,IMMT1,IMMT2: NG
  AS_WFS,POP_POM,POS_POM,REFL_WFS,POM1: NG
   => Confirmation only, no recovery work done.

 OK:
  ETMX,ETMY,ITMX,ITMY
  BS(IM,BF),SRM(IM,BF),SR2(IM,BF,STM),SR3(IM,BF)
  PRM(IM,BF),PR2(IM,BF),PR3(IM,BF)
  MCF(MCO,MCI),MCE
  OMMT1,OMMT2,OSTM

  PCAL(EX1,EX2,EY1,EY2):Not confirmed because it is ON only when used.

Temperature Control
 [Checked]
 Check if the temperature acquisition is working properly.

 [Check result]
  No problem
 

Non-image files attached to this comment
satoru.ikeda - 17:39 Friday 22 October 2021 (18669) Print this report

Routine check for DGS maintenance day (Oct. 22)

Stepper motor
 [Checked]
  Check if the power supply can be controlled by BIO.
  Check if the script can be started.
  Confirmed that the limit switch installed is in Green.
 [Check result]
  No problem

Release Sensor
 [Checked]
 Confirmed that the Release Sensor is 0. 1: Seated
 [Check result]
 ETMY, ITMX, ITMY : No problem
 ETMX_F1_1 : abnormal, no other problems (known)

Picomotor
 [Checked]
  PingTest is performed.
 [Result]
 NG:
  No problem

  MCI,STM1,IMMT1,IMMT2: NG
  AS_WFS,POP_POM,POS_POM,REFL_WFS,POM1: NG
   => Confirmation only, no recovery work done.

 OK:
  ETMX,ETMY,ITMX,ITMY
  BS(IM,BF),SRM(IM,BF),SR2(IM,BF,STM),SR3(IM,BF)
  PRM(IM,BF),PR2(IM,BF),PR3(IM,BF)
  MCF(MCO,MCI),MCE
  OMMT1,OMMT2,OSTM

  PCAL(EX1,EX2,EY1,EY2):Not confirmed because it is ON only when used.

Temperature regulator
 [Checked]
 Check if the temperature acquisition is working properly.

 [Check result]
  No problem

Images attached to this comment
Non-image files attached to this comment
Search Help
×

Warning

×